# EMULATE: A Multi-Agent Framework for Determining the Veracity of Atomic Claims by Emulating Human Actions

### v1

@@ Abstract
Spencer Hong, Meng Luo, Xinyi Wan National University of Singapore
@@ 1 Introduction
Figure 1: Claim veracity classification algorithm. The input to the algorithm is an atomic claim c along with two values specifying the maximum number of search queries that can be made and the maximum number of search results returned per query. The output of the algorithm is a binary label indicating the veracity of the claim. Each LLM agent is highlighted in yellow.
@@ 4 Experiments
Baselines. We compare our multi-agent system with four baselines: (1) FacTool (Chern et al., 2023), (2) FactCheck-GPT (Wang et al., 2024), (3) SAFE (Wei et al., ), and (4) FIRE Wei et al. . Note that FIRE (Xie et al., 2025) is the only baseline that was designed to take as input an atomic claim and output True or False (like EMULATE). In each of the other three baselines, checking the veracity of atomic claims is one step in the algorithm, which means that modifications are required to make comparisons.
Implementation. For our main experiments, we employ OpenAI’s GPT-4.1 model111gpt-4.1-2025-04-14 with a temperature of 1 for all agents in EMULATE as well as the baseline systems. All EMULATE agents are provided with zero-shot prompts that contain instructions for the subtasks. Unless otherwise stated, MAX_SEARCH_QUERIES and MAX_SEARCH_RESULTS_PER_QUERY are set to 4 and 2 respectively. To make web searches, we invoke API calls with serper.dev.
@@ 5 Results
Lastly, we run experiments on the entire GPT-4.1 model family to determine if EMULATE still works well when the underlying LLM of each agent is supplanted with a weaker model. According to Figure 2, as the underlying LLM weakens, the weighted-F1 scores decrease as well. Intuitively, weaker models are expected to be less performant on the subtasks in EMULATE, which can lead to suboptimal results. However, we can see that the performance of EMULATE when equipped with GPT-4.1-mini is sometimes close to the performance with GPT-4.1.

### v2

@@ Abstract
Spencer Hong  Meng Luo  Xinyi Wan National University of Singapore {spencer.hong, mluo, wan.xinyi}@u.nus.edu
@@ 1 Introduction
Figure 1: Claim veracity classification algorithm. The input to the algorithm is an atomic claim c along with two values specifying the maximum number of search queries that can be made and the maximum number of search results returned per query. The output of the algorithm is a binary label indicating the veracity of the claim. Each LLM agent is highlighted in green.
@@ 4 Experiments
Baselines. We compare our multi-agent system with four baselines: (1) FacTool (Chern et al., 2023), (2) FactCheck-GPT (Wang et al., 2024), (3) SAFE (Wei et al., 2024), and (4) FIRE (Xie et al., 2025). Note that FIRE (Xie et al., 2025) is the only baseline that was designed to take as input an atomic claim and output True or False (like EMULATE). In each of the other three baselines, checking the veracity of atomic claims is one step in the algorithm, which means that minor modifications to the corresponding open-source repositories were required to make comparisons [^1] For FactCheck-GPT, we also modify the code to utilize serper.dev to obtain a maximum of 10 URLs per query..
Implementation. For our main experiments, we employ OpenAI’s GPT-4.1 model222gpt-4.1-2025-04-14 with a temperature of 1 for all agents in EMULATE as well as the baseline systems. All EMULATE agents are provided with zero-shot prompts that contain instructions for the subtasks. Unless otherwise stated, MAX_SEARCH_QUERIES and MAX_SEARCH_RESULTS_PER_QUERY are set to 4 and 2 respectively. To make web searches, we invoke API calls with serper.dev.
@@ 5 Results
Lastly, we run experiments on the entire GPT-4.1 model family to determine if EMULATE still works well when the underlying LLM of each agent is supplanted with a weaker model. According to Figure 2, as the underlying LLM weakens, the weighted-F1 scores decrease as well. Intuitively, weaker models are expected to be less performant on the subtasks in EMULATE, which can lead to suboptimal results; however, we can see that the performance of EMULATE when equipped with GPT-4.1-mini is sometimes close to the performance with GPT-4.1.